using substution mehhod [x+2y=11]
[2x-y=2

Answer by algebrahouse.com(1659)

x + 2y = 11 -----> x = -2y + 11 {subtracted 2y from both sides}
2x - y = 2

2x - y = 2 {bottom equation}
2(-2y + 11} - y = 2 {substituted -2y + 11, in for x, into bottom equation}
-4y + 22 - y = 2 {used distributive property}
-5y + 22 = 2 {combined like terms}
-5y = -20 {subtracted 22 from both sides}
y = 4 {divided both sides by -5}

x = -2y + 11 {re-arranged top equation}
x = -2(4) + 11 {substituted 4, in for y, into top equation}
x = -8 + 11 {multiplied -2 by 4}
x = 3 {combined like terms}

x = 3 and y = 4

For more help from me, visit: www.algebrahouse.com